OpenStudy (anonymous):

4. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

Please help I don't understand

OpenStudy (anonymous):

0.01 is the accepted tolerance for the drill bit

OpenStudy (anonymous):

There's supposed to be two answers to the equation.

OpenStudy (anonymous):

if there is a tolerance that means the drill bit can be -slightly above or -slightly below the specified diameter.

OpenStudy (anonymous):

The first equation is x - 0.0525 = 0.01. The 2nd is: x - 0.0525 = -0.01

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Is the answer to the first one 0.0625?

OpenStudy (anonymous):

0.0625 - 0.0525 = 0.1, right?

OpenStudy (anonymous):

Right, and the answers to both of those equations will equal the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the second diameter (minimum diameter) gives a negative number in the absolute bracket, and the absolute again gives the required 0.1

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so the second one is x=0.0425 0.0425 - 0.0525 = -0.01, |-0.01|=0.01

OpenStudy (anonymous):

like already stated above

OpenStudy (anonymous):

So the answer to the second one is 0.0425, correct?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

minimum diameter is 0.0425

OpenStudy (anonymous):

So the minimum diameter of the drill bits is 0.425. And the maximum is 0.625.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

both have a deviation of no more than 0.01

OpenStudy (anonymous):

Thanks a lot.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Could you help me on another? But slightly different?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x + 6/8 = 7x – 1/6 24/1(3x + 6/8) = 24/1(7x – 1/6) 3(3x + 6) = 4(7x – 1) 9x + 18 = 28x – 1) 9x – 9x + 18 = 28x – 9x – 1 18 = 19x – 1 18 + 1 = 19x – 1 + 1 19 = 19x 1 = x

OpenStudy (anonymous):

@mathessentials

OpenStudy (anonymous):

he could substitute the x in the original equation..

OpenStudy (anonymous):

I figured it out on my own but I seemed to have gotten it wrong too

OpenStudy (anonymous):

Here's my "revised" version: The mistake he made was on Line #5. When he multiplied 4 x 7x he didn’t multiply 4 x 1 which was also in the parenthesis. Here’s my solution: 3x + 6/8 = 7x – 1/6 24/1(3x + 6/8) = 24/1(7x – 1/6) 3(3x + 6) = 4(7x – 1) 9x + 18 = 28x – 4 9x – 9x + 18 = 28x – 9x – 4 18 = 19x – 4 18 + 4 = 19x – 4 + 4 22 = 19x 1.2 ≈ x

OpenStudy (anonymous):

But I checked my answer and it's not right.

OpenStudy (anonymous):

the fist difference is in the line 9x – 9x + 18 = 28x – 9x – 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

one before actually

OpenStudy (anonymous):

That's what I said.

OpenStudy (anonymous):

yes, max did not multiply the second part of the paranthesis

OpenStudy (anonymous):

He did do 4 x 7x but didn't do 4 x 1...

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Yes, but what did I do wrong? ANd did Max do anything else wrong before or after that?

OpenStudy (anonymous):

3x + 6/8 = 7x – 1/6 24/1(3x + 6/8) = 24/1(7x – 1/6) what did he try to do there?

OpenStudy (anonymous):

I'm not sure I don't know where he got the "24/1" from

OpenStudy (anonymous):

yeah.... thats hard to understand

OpenStudy (anonymous):

This is what the equation originally looked like:

OpenStudy (anonymous):

I had to copy & paste it into a Word Document.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Np just need to get this done

OpenStudy (anonymous):

the denominators in the brackets are 6x8, which = 48, probably has something to do with the 24

OpenStudy (anonymous):

he tries to get the same common denominator

OpenStudy (anonymous):

Oh yeah he broke 48 down...

OpenStudy (anonymous):

what he should have done \[\frac{ 6 \times (3x+6) }{6 \times 8 }=\frac{ 8 \times (7x-1) }{8 \times 6 }\]

OpenStudy (anonymous):

Oh he should have multiplied the denominators?

OpenStudy (anonymous):

Its clear

OpenStudy (anonymous):

he should get the same denominators by multiplying the one with the other and multiplying the numerator with the "foreign" denominator because you're allowed to multiply by any fraction that makes one, like 8/8 or 7/7 or 6/6, when you multiply with a number down must also with the same one up

OpenStudy (anonymous):

So could you figure out the problem? And walk me through it?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } + \frac{ 2 }{ 5 } =\frac{ 5 \times 1 }{ 5 \times 3 } + \frac{ 3 \times 2 }{ 3 \times5 }\]

OpenStudy (anonymous):

Where did you get those fractions from?

OpenStudy (anonymous):

they are example!!

OpenStudy (anonymous):

Ohhhhhh

OpenStudy (anonymous):

on how to get a common denominator it's the same principle in the equation

OpenStudy (anonymous):

Modulus is given so for min. answer modulus will be negative i.e., |x − 0.0525| will be --(x-0.0525) and for max it will be +(x-0.0525) then solve to get the the value of x in each case

OpenStudy (anonymous):

So is this what I have so far? \[\frac{ 3x + 6 }{ 40 } = \frac{ 7x - 1 }{ 40 }\]

OpenStudy (anonymous):

8x6 is 48

OpenStudy (anonymous):

@Addu5221 we're on a different question.

OpenStudy (anonymous):

My bad...

OpenStudy (anonymous):

ohk sorry

OpenStudy (anonymous):

and also you must do the operations on both parts of the fraction so also make 6(3x + 4)

OpenStudy (anonymous):

3x+6*

OpenStudy (anonymous):

the fraction itself should not be changed in value so if multiply below must multiply above

OpenStudy (anonymous):

This is what we have so far right? \[\frac{ 3x + 6 }{ 48 } = \frac{ 7x - 1 }{ 48 }\]

OpenStudy (anonymous):

\[\frac{ 6 \times (3x+6) }{ 48 }=\frac{ 8 \times (7x-1) }{ 48 }\]

OpenStudy (anonymous):

Oh, okay so we have to do: 6 x 3x + 6 = 8 x 7x - 1?

OpenStudy (anonymous):

yes see above

OpenStudy (anonymous):

Hold on let me type this out on my Word Document

OpenStudy (anonymous):

But revised version is perfect

OpenStudy (anonymous):

if you changed only the denominator, and the numerator is the same then: both (x + ) terms are divided by the same number, which is not what the problem stated. they have different fractions and if we want them to have the same fractions we must compensate that

OpenStudy (anonymous):

Okay, I'm back.

OpenStudy (anonymous):

what is this discussion for greenlegodude your revise version is correct

OpenStudy (anonymous):

@Addu5221 This is the question we are working on: Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x + 6/8 = 7x – 1/6 24/1(3x + 6/8) = 24/1(7x – 1/6) 3(3x + 6) = 4(7x – 1) 9x + 18 = 28x – 1) 9x – 9x + 18 = 28x – 9x – 1 18 = 19x – 1 18 + 1 = 19x – 1 + 1 19 = 19x 1 = x

OpenStudy (anonymous):

Yes I get that Mathessentials

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

so where are we at

OpenStudy (anonymous):

So 6 x 3x + 6. We solve this first. So: subtract 6 from both sides?

OpenStudy (anonymous):

@Addu5221 Here is the exact mistake Max made:

OpenStudy (anonymous):

So 6 x 3x + 6. We solve this first. So: subtract 6 from both sides? So then we have 0 x 3x? I don't get it....

OpenStudy (anonymous):

"6 x 3x + 6." where from?

OpenStudy (anonymous):

Um, that's from the numerator of the first fraction.

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

OK so what max did was this: he found the smallest number that is divisible by 8 and by 6

OpenStudy (anonymous):

Since we multiplied 6 to the denominator of the 1st fraction. So we have to do it to the numerator, correct?

OpenStudy (anonymous):

RIght

OpenStudy (anonymous):

yes the numerator would be 48

OpenStudy (anonymous):

You mean the denominator right?

OpenStudy (anonymous):

you can always get a common denominator by multiplying the different ones together however there's also a smaller multiplication sometimes that gives the same result

OpenStudy (anonymous):

you are right, the denominator

OpenStudy (anonymous):

e.g. 8 and 6 48 is a common one however 24 also is (3 and 4)

OpenStudy (anonymous):

max chose the smallest possible whihc is 24 instead of 48 that is not an error

OpenStudy (anonymous):

the first error he did was by skipping 4x -1

OpenStudy (anonymous):

Yes I know about getting a common denominator like say if youi have 1/2 and 3/4 the common denominator is 4 because 2 x 2 = 4 and 4 x 1 = 4. It's not 8 because there is a smaller number that can go into both of these.